OpenStudy (rebecca1588):

What is the y-intercept of the line? y=-7x+1 y-intercept =?

OpenStudy (anonymous):

the y intercept is 1

OpenStudy (anonymous):

y=mx+b b=y intercept

OpenStudy (anonymous):

When x=0---> y=-7(0)+1 Y=1

OpenStudy (rebecca1588):

Thanks everyone! :)